uiz Instructions Question 2 3 pts A sample data set with a bell-shaped distribution has mean2 and standard deviation s 1.1. Find the approximate proportion of observations in the data set that lie: 1. below -0.2; [Select 2. below 3.1; I Select ] 3. between -1.3 and 0.9. [ Select ] Previous Next
Answer #1

1)

for normal distribution z score =(X-μ)/σx
here mean=       μ= 2
std deviation   =σ= 1.1000
probability = P(X<-0.2) = P(Z<-2)= 0.0228

2)

probability = P(X<3.1) = P(Z<1)= 0.8413

3)

probability = P(-1.3<X<0.9) = P(-3<Z<-1)= 0.1587-0.0013= 0.1574